OpenStudy (anonymous):

A skirt costs $20 more than the blouse. The skirt costs $65. What is the price of the blouse. Write an equation. Then solve the problem. a. $ 85 c. $55 b. $ 45 d. $ 35

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ummm....not A Think. Ignore their instructions to write an equation. What is the answer. Use common sense

OpenStudy (anonymous):

waht is it please

OpenStudy (anonymous):

Oh for crying out loud! You can't be serious! A skirt costs $20 more than the blouse. The skirt costs $65. How much is the blouse????????????????

OpenStudy (anonymous):

A beer costs 2 dollar more than a coke. A beer is $4. How much is a Coke?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

2

OpenStudy (anonymous):

6 is 2 dollars more than a beer!! I said a beer is 2 dollars more than the Coke.

OpenStudy (anonymous):

If five pounds of potatoes cost $3.45, then 2 lb cost $.89.

OpenStudy (anonymous):

t or f

OpenStudy (anonymous):

Horrible math teachers have really led to people giving up on actually thinking. Always looking for the right FORMULA or PROCEDURE....and not just thinking about things a tiny bit. Depressing

OpenStudy (anonymous):

Which equation represents “a number increased by 8 is equal to sixteen”? A. y – 8 = 16 B. y + 8 = 16 C. 8 – y = –16 D. 8 + y = –16

OpenStudy (anonymous):

a number increased by 8 is equal to sixteen x +8 = 16

OpenStudy (anonymous):

You are saving $5 each week to attend a summer program that costs $300. You already have $84. You have also received a $50 scholarship for the program. How many weeks of saving do you need? a. 44 weeks c. 27 weeks b. 50 weeks d. 34 weeks

OpenStudy (texaschic101):

s = b + 20 (skirt costs $20 more then a blouse) s = 65 now sub 65 in for s 65 = b + 20 -- subtract 20 from both sides 65 - 20 = b 45 = b thats your first one... ===================== 84 + 50 + 5x = 300 (x stands for # of weeks) 134 + 5x = 300 -- subtract 134 from both sides 5x = 300 - 134 5x = 166 x = 166/5 x = 33.2 weeks My answer is not an answer choice....closest is 34 weeks.